# NavCoT: Boosting LLM-Based Vision-and-Language Navigation via Learning Disentangled Reasoning

### v1

@@ E.2.3 Generalization on Other Datasets
To further verify the generalization of NavCoT on different VLN datasets with much longer instructions (RxR [5]) and high-level instructions (REVERIE [2]), we conduct experiments on RxR and REVERIE, and the results are given in Table VI and Table VII, respectively. Since there are no reported results of language-only backbone on RxR and REVERIE, we develop a naive baseline, which is a variant of NavCoT that makes direct action predictions (DAP) rather than generating navigational chain-of-thoughts. We also present the results of a strong pretrained cross-modal backbone HAMT [30] with the same training strategy as ours. From Table VI we can find that although NavCoT shows the performance gap with the cross-modal methods, it significantly outperforms DAP especially in the CLS and nDTW metrics, showing that our method not only improves the navigation accuracy but also enables better instruction following, which is crucial for long-horizon navigation. Table VII shows that different methods encounter significant performance drop on REVERIE compared to R2R, revealing the challenge of navigating under high-level instructions with limited information. However, NavCoT is still superior over DAP in a large margin, showing the effectiveness of the proposed method.
@@
TABLE VIII: Low-resource experimental results on Val Unseen on R4R and R2R. * denotes our re-implementation results with the imitation learning strategy [30]. In R4R, CLS and NDTW are the main metrics.
@@
Figure 6: Action decision visualization of NavCoT. The ground-truth actions are annotated by red boxes.
@@ E.2.4 Low-Resource Experiments
Table VIII shows that under the low-resource setting with the same training strategy, the performance gap between NavCoT and HAMT [30] is largely reduced compared to that in Table I. For example, the performance gap of CLS between HAMT [30] and NavCoT is reduced to $\sim$8. Moreover, NavCoT outperforms DAP in a large margin on both R2R and R4R, showing the good generalization ability of NavCoT under low-resource settings. Note that the results on R4R show that NavCoT surpasses DAP significantly in OSR, CLS, and NDTW, demonstrating that beyond correct action decisions, NavCoT also shows promising instruction following ability for long trajectories.
@@
Figure 8: Failure cases. The ground-truth (GT) actions and the actions chosen by NavCoT are denoted by green and red boxes, respectively.
@@ E.3 Visualization
Action decision visualizations. Fig. 6 gives the action decision visualization of NavCoT, where we can find that NavCoT generates reasonable navigational reasoning for guiding itself to make action decisions. For instance, in Step 2, from the history and observations, NavCoT infers that its position is possibly the hall and the option C may contain the desired table, then outputs the imagination that leads to the correct action. Notably, due to the domain gap between NavCoT and VLM, the observations inevitably do not always contain the imagination (Step 1, 3 and 4). However, with in-domain training, NavCoT shows the emerging ability to recognize potential connection between observations and imaginations. For example, in Step 4, the imagination office does not directly match any observation, whereas NavCoT correctly chooses the option C containing a wooden desk which suggests a scene of office.
@@
Failure cases. We further analyze some failure cases of NavCoT in Fig. 8. We find two typical types of failures: 1) Indistinguishable observation descriptions. As shown in Fig. 8(a), since the observations in Options C and D in Step 2 are visually similar, their textual descriptions are also indistinguishable. Such cases also impose a big challenge on the cross-modal navigation agent. 2) Loss of global observation information. In Fig. 8(b), instruction mentions the first door on the left, which can not be easily inferred from independent candidate observation descriptions. The introduction of panoramic observation information may further help improve the global scene understanding.

### v2

@@ E.2.3 Generalization on Other Datasets
To further verify the generalization of NavCoT on different VLN datasets with much longer instructions (RxR [5]) and high-level instructions (REVERIE [2]), we conduct experiments on RxR and REVERIE, and the results are given in Table IV and Table V, respectively. Since there are no reported results of language-only backbone on RxR and REVERIE, we develop a naive baseline, which is a variant of NavCoT that makes direct action predictions (DAP) rather than generating navigational chain-of-thoughts. We also present the results of a strong pretrained cross-modal backbone HAMT [34] with the same training strategy as ours. From Table IV we can find that although NavCoT shows the performance gap with the cross-modal methods, it significantly outperforms DAP especially in the CLS and nDTW metrics, showing that our method not only improves the navigation accuracy but also enables better instruction following, which is crucial for long-horizon navigation. Table V shows that different methods encounter significant performance drop on REVERIE compared to R2R, revealing the challenge of navigating under high-level instructions with limited information. However, NavCoT is still superior over DAP in a large margin, showing the effectiveness of the proposed method.
@@
TABLE VI: Low-resource experimental results on Val Unseen on R4R and R2R. * denotes our re-implementation results with the imitation learning strategy [34]. In R4R, CLS and NDTW are the main metrics.
@@
Figure 6: Landmark extraction visualization of different datasets.
Figure 7: Action decision visualization of NavCoT. We only extract two steps for illustration and give the whole trajectory in the supplementary material. The ground-truth actions are annotated by red boxes.
@@ E.2.4 Low-Resource Experiments
Table VI shows that under the low-resource setting with the same training strategy, the performance gap between NavCoT and HAMT [34] is largely reduced compared to that in Table I. For example, the performance gap of CLS between HAMT [34] and NavCoT is reduced to $\sim$8. Moreover, NavCoT outperforms DAP in a large margin on both R2R and R4R, showing the good generalization ability of NavCoT under low-resource settings. Note that the results on R4R show that NavCoT surpasses DAP significantly in OSR, CLS, and NDTW, demonstrating that beyond correct action decisions, NavCoT also shows promising instruction following ability for long trajectories.
Figure 8: Reasoning output visualization of different Chain-of-Thought (CoT) methods. Apart from NavCoT (ours), all other comparison methods are zero-shot.
@@ E.3 Visualization
Since we utilize the LLM for instruction landmark extraction to collect imagination GTs, we further validate the accuracy of landmark extraction. Specifically, we manually check the landmark extraction results of our used LLM, i.e., ChatGPT [61] for the R2R, REVERIE and RxR datasets. For each dataset, we randomly select 100 instructions for evaluation. The landmark extraction is considered successful if every landmark in it is extracted. We find that almost all landmarks can be successfully extracted by the ChatGPT, and the success rate of landmark extraction is 100% for R2R, 98% for REVERIE, and 98% for RxR. Since the landmarks of an R4R instruction can be combined from the landmarks of the corresponding R2R instructions, the landmark extraction accuracy of R4R instructions is consistent with that of R2R instructions. We also present the visualization results of landmark extraction for different datasets in Fig. 6, where we can observe that ChatGPT extracts landmarks accurately for instructions of different datasets, demonstrating that it is robust to instructions with different styles. The accurate landmark extraction effectively ensures the correctness of our subsequent collection of the Imagination GTs.
Action decision visualizations. Fig. 7 gives the action decision visualization of NavCoT, where we can find that NavCoT generates reasonable navigational reasoning for guiding itself to make action decisions. For instance, in Step 2, from the history and observations, NavCoT infers that its position is possibly the hall and the option C may contain the desired table, then outputs the imagination that leads to the correct action. Notably, due to the domain gap between NavCoT and VLM, the observations inevitably do not always contain the imagination. However, with in-domain training, NavCoT shows the emerging ability to recognize potential connection between observations and imaginations. For example, in Step 4, the imagination office does not directly match any observation, whereas NavCoT correctly chooses the option C containing a wooden desk which suggests a scene of office.
Comparison of different CoT methods. Figure 8 presents the visualization comparison of our NavCoT with other three zero-shot CoT methods: the original CoT (GPT-4)*, the original CoT (LLaMA2)*, and the NavCoT (LLaMA2)*. From Figure 8, we can observe two obvious advantages of our NavCoT over the other CoT approaches: 1) Through formalizing navigational reasoning into three key steps: future imagination, visual information filtering, and action decision, NavCoT significantly reduces the redundant information caused by the domain gap between the LLM training corpus and VLN task in its reasoning output. In contrast, the reasoning of the original CoT (GPT-4)* contains extensive redundancy, and the reasoning of the original CoT (LLaMA2)* contains a large amount of noise. For instance, the original CoT (GPT-4)* mixes the instruction, common sense, and history in its output, and produces redundant reasoning which is hard to extract useful information from it. The original CoT (LLaMA2)* outputs the wrong reasoning that “the history contains the bedroom” and its chosen action D (the direction is go back) does not match its reasoning that “the instruction asks me to veer left”. 2) By introducing the parameter-efficient in-domain training strategies, NavCoT can produce reasonable formalized imagination and uses the imagination to guide itself for choosing the correct action decision. For example, our NavCoT produces proper imagination “stairs” and matches the imagination to the correct option B which contains the stairs for successful navigation. In contrast, the NavCoT (LLaMA2)* fails to match its imagination output “go down the stairs” to the correct option B and instead chooses option D.
@@ F Conclusion
Acknowledgment
This work was supported by National Key Research and Development Program of China (2024YFE0203100), National Natural Science Foundation of China (NSFC) under Grants No.62476293, Shenzhen Science and Technology Program (Grant No.GJHZ20220913142600001), Nansha Key R&D Program under Grant No.2022ZD014, and General Embodied AI Center of Sun Yat-sen University.